A certain recipe asks for 4/5 cup of granola how much is needed to make 2/9 of a recipe
Answer by stanbon(75887)

A certain recipe asks for 4/5 cup of granola
how much is needed to make 2/9 of a recipe
-------
Ans: (4/5)(2/9) = 8/45 = 0.1778 cup
-------------
Cheers,
Stan H.
------------